The draft statement of financial position of Linfield Limited for the year ended 31 December 2016 are as below: Statement of financial position as at 31 December 2016 (with comparative figures) 2016 2015 Non-current assets Property, plant and equipment 35,750 281,000 Investment properties 7,000 7,500 Intangibles assets 6,000 6,200 49,250 294,200 Current assets Inventories 3,980 4,545 Trade receivables 7,410 9,075 Short-term investments 500 Cash and bank 7,555 13,650 26,105 20,610 14,810 375,355 Total assets Equity and reserves Share capital 54,500 48,800 Retained profits 190,000 39,945 244,500 188,745 Current liabilities Trade payables Other payables 8,745 9,340 1,100 800 Interest payables 2,900 3,000 Tax payables 11,725 10,125 Bank overdraft 10,500 24,470 33,765
2016 2015 Non-current liabilities Bank loans 92,500 82,500 Deferred tax liabilities 13,885 9,800 106,385 92,300 375,355 14,810 Total equity and liabilities The following information is available: 1. Profits before tax for the year $98,522 million has been arrived at after charging (crediting) the following items: Finance cost 6,540 Depreciation charge 4,798 Change in fair value of investment properties Gain on sale of property, plant and equipment (500) (80) Inventories written down 18 Provision for bad debts Impairment loss of intangible assets 200 2. In March 2016, Linfield Limited sold a machine for $500 million. The c arrying amount of the machine in the books was $420 million. 3. The short-term investments represented marketable securities was purchased on 29 December 2016. I matured on 25 March 2017 There were new issues of ordinary shares at market price during the year. Dividends paid 4. during the year amounted to $38,127 million. 5. It was Linfield Limited's accounting policy to perform impairment loss test on its assets at each financial year end. Impairment losses, if any, were written off immediately as expenses 6. Income tax expense charged to the statement of profit or loss was $10,340 million.
Required: (a) Prepare the statement of cash flows for the year ended 31 December 2016, using the indirect method to determine the cash flows from operating activities, for Linfield Limited (b) Which companies are required to prepare interim reports and why? Briefly discuss the purpose of interim reporting.
Answer #1
Cash Flow Statement
Indirect Method
Cash Flow from Operating Activities
Net Income $   98,522.00
Adjustments
Finance Cost $      6,540.00
Depreciation $      4,798.00
Change in Fair Value of Investment $       (500.00)
Gain on Sale of PPE $         (80.00)
Inventories written down $            18.00
Provision for bad debts $              5.00
Impairment loss of intangible assets $         200.00
Changes in Current Assets/ Current Liabilities
Increase in Inventories $       (583.00) =3980-4545-18
Decrease in Trade Receivables $      1,660.00 =9075-7410-5
Increase in Short term Investments $       (500.00)
Decrease in Trade Payables $       (595.00) =8745-9340
Increase in Other Payables $         300.00 =1100-800
Increase in Tax Payables $      1,600.00 =11725-10125
Less Income Tax paid $   (6,255.00) =-(10340-13885+9800)
Total Adjustments $      6,608.00
Cash from Operating Activities $ 105,130.00
Cash flow from Investing Activities
Sale of Equipment $         500.00
Purchase of PPE $ (59,968.00) =281000-420-4798-335750
Net cash used In investing activities $ (59,468.00)
Cash flow from Financing Activities
Issue of Share Capital $      5,700.00 =54500-48800
Borrowing of loan from bank $   10,000.00 =92500-82500
Finance Cost $   (6,640.00) =-(6540+3000-2900)
Dividend Paid $ (38,127.00)
Net cash used in financing activities $ (29,067.00)
Increase in Cash $    16,595.00
Opening Balance of Cash & Equivalents $    (2,945.00) =7555-10500
Ending Balance of Cash & Equivalents $    13,650.00

Bank Overdraft has been considered as part of cash.

(B) It is required by the companies which are publicaly held because investor would like to know how the company is performing throughout the year so that they can take informed decisions, it is also helpful for government authorities to check on their tax compliance.

Purpose of interim reporting is to
To estimate the earnings of the company
To make projections
To evaluate management performance regularly